OpenStudy (anonymous):

evaluate (1/2)^-3 x (3^-2)

OpenStudy (anonymous):

8/9

OpenStudy (anonymous):

i'll help

OpenStudy (anonymous):

(1/2)^-3 =8 3^-2=1/9 8*1/9 =8/9

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2 })^{-3}\times(3^{-2}) = (2)^{3} \times (\frac{ 1 }{3^{2} })\]